# UAV-Mounted IRS (UMI) in the Presence of Hovering Fluctuations: 3D Pattern Characterization and Performance Analysis

### v2

@@ IV-C SISO-Active-IRS
$F_{\gamma_{max}|\mathbb{G}}(\gamma_{\text{th}})$ $=F_{U}(\mu_{w}(\mathbb{G}))$ (75)
$+F_{U}^{(1)}(\mu_{w}(\mathbb{G}))\mathbb{E}_{W}\{(w(\mathbb{G})-\mu_{w}(\mathbb{G}))\}$
$+F_{U}^{(2)}(\mu_{w}(\mathbb{G}))\mathbb{E}_{W}\{(w(\mathbb{G})-\mu_{w}(\mathbb{G}))^{2}\}.$
Given that $\mathbb{E}_{W}\{(w(\mathbb{G})-\mu_{w}(\mathbb{G}))\}=0$, we can rewrite (75) as
$F_{\gamma_{max}|\mathbb{G}}(\gamma_{\text{th}})=Q(\frac{\mu_{v}-\sqrt{\mu_{w}(\mathbb{G})}}{\sigma_{v}})+F_{U}^{(2)}(\mu_{w}(\mathbb{G}))\sigma_{w}^{2}(\mathbb{G}),$ (76)
where
$F_{U}^{(2)}(\mu_{w}(\mathbb{G}))$ $=\frac{1}{{4\sqrt{2\pi}\sigma_{v}^{3}}}\exp(\frac{-(\mu_{v}-\sqrt{\mu_{w}(\mathbb{G})})^{2}}{2\sigma^{2}_{v}})$ (77)
×$\times({\frac{\mu_{v}}{\mu_{w}(\mathbb{G})}-\frac{1}{(\mu_{w}(\mathbb{G}))^{\frac{1}{2}}}-\frac{\sigma_{v}^{2}}{(\mu_{w}(\mathbb{G}))^{\frac{3}{2}}}}).$
Finally, by substituting (76) into (70), the OP is obtained. ∎
IV-D MISO-Active-IRS
In this section similar to the MISO-Passive-IRS section, we consider only the LOS part in the MISO system. Hence, the optimal active-IRS reflection design in (62) and (63) can be rewritten as
$[\mathbf{\Phi}^{*}]_{n}=e^{-j\angle[\bar{\mathbf{h}}_{IU}^{H}]_{n}},~{}\forall n,$ (78)
$(A^{*})^{2}=\frac{P_{F}}{P_{t}\|\mathbf{H}_{BI}\|^{2}+NN_{F}},$ (79)
where $\|\mathbf{H}_{BI}\|^{2}=MN\beta_{0}$. Therefore, by substituting (78) and (79) into (12), the maximum instantaneous SNR of MISO-Active-IRS system is obtained as
$\gamma_{\text{max}}=\frac{(\sqrt{\beta_{0}}\sum_{n=1}^{N}{|\bar{h}_{IU,n}|})^{2}}{c_{4}\underbrace{\sum_{n=1}^{N}{|\bar{h}_{IU,n}|^{2}}}_{Z_{1}}+c_{5}/\mathbb{G}},$ (80)
where $c_{4}=\frac{N_{F}}{P_{t}M}$, and $c_{5}=\frac{NN_{0}}{P_{F}}({\beta_{0}}+\frac{N_{F}}{P_{t}M})$.
Proposition 3.
In the case of CLT approximation, the OP of the MISO-Active-IRS system can be obtained as
$P_{\text{out}}=\sum_{i=1}^{J}p(i)R_{2}(q(i)),$ (81)
where
$R_{2}(\mathbb{G})=Q(\frac{\mu_{v}-\sqrt{\mu_{\varpi}(\mathbb{G})}}{\sigma_{v}})+F_{U}^{(2)}(\mu_{\varpi}(\mathbb{G}))\sigma_{\varpi}^{2},$ (82)
$F_{U}^{(2)}(\mu_{\varpi}(\mathbb{G}))$ $=\frac{1}{{4\sqrt{2\pi}\sigma_{v}^{3}}}\exp(\frac{-(\mu_{v}-\sqrt{\mu_{\varpi}(\mathbb{G})})^{2}}{2\sigma^{2}_{v}})$ (83)
×$\times({\frac{\mu_{v}}{\mu_{\varpi}(\mathbb{G})}-\frac{1}{(\mu_{\varpi}(\mathbb{G}))^{\frac{1}{2}}}-\frac{\sigma_{v}^{2}}{(\mu_{\varpi}(\mathbb{G}))^{\frac{3}{2}}}}).$
Here, $\mu_{\varpi}(\mathbb{G})=c_{4}\gamma_{th}\mu_{Z_{1}}+\frac{c_{5}\gamma_{th}}{\mathbb{G}}$ and $\sigma^{2}_{\varpi}=\gamma_{th}^{2}c^{2}_{4}\sigma^{2}_{Z_{1}}$.
Proof.
We can rewrite SNR in (80) as
$\gamma_{max}=\frac{U_{2}}{c_{4}Z_{1}+c_{5}/\mathbb{G}},$ (84)
where $V_{2}=\sqrt{\beta_{0}}\sum_{n=1}^{N}{|\bar{h}_{IU,n}|}$ and $U_{2}=V_{2}^{2}$. Therefore, the OP can be given as
[…]
where $f_{\mathbb{G}}(\mathbb{G})$ is given by (39). $F_{\gamma_{max}|\mathbb{G}}(.)$ is the CDF of $\gamma_{\text{max}}$ conditioned on the radiation gain, i.e., $\mathbb{G}$, which can be obtained similar to steps taken in deriving (71) outlined in Theorem 4 as follows
$F_{\gamma_{max}|\mathbb{G}}(\gamma_{\text{th}})=\mathbb{E}_{\varpi}\{F_{U_{2}}(\varpi)\}$ (86)
where $\varpi=c_{4}\gamma_{th}Z_{1}+c_{5}\gamma_{th}/\mathbb{G}$. The distribution of $\varpi$ follows $\mathcal{N}(\mu_{\varpi}(\mathbb{G}),\sigma_{\varpi}^{2}(\mathbb{G}))$ where $\mu_{\varpi}(\mathbb{G})=c_{4}\gamma_{th}\mu_{Z_{1}}+\frac{c_{5}\gamma_{th}}{\mathbb{G}}$ and $\sigma^{2}_{\varpi}=\gamma_{th}^{2}c^{2}_{4}\sigma^{2}_{Z_{1}}$. Note that $F_{\gamma_{max}|\mathbb{G}}$ is calculated same as (76) by replacing variable $v$ and $w$ with $v_{2}$ and $\varpi$, respectively. By substituting $F_{\gamma_{max}|\mathbb{G}}$ into (85), the OP is derived. ∎
@@ V Numerical results and discussions
This section provides numerical results to validate the theoretical analysis of the UMI system. The simulation parameters are configured as follows: the BS, IRS, and UE are located at $[0,0,20]$, $[10,10,70]$ and $[50,50,0]$, respectively. The path-loss exponents for BS-IRS and IRS-UE links are configured as $\alpha_{1}=2$ and $\alpha_{2}=2.2$, respectively, with a reference path-loss of $c_{0}=-30$ dB applied to all links [34]. Additionally, The Rician factors for the BS-IRS and IRS-UE channels are set to $K_{0}=10$ dB and $K_{1}=5$ dB, respectively. The thermal noise power of the user’s receiver and active IRS is $N_{0}=-80$ dBm and $N_{F}=-70$ dBm, respectively [31]. The active IRS amplification power $P_{F}=0.1P_{t}$ and SNR threshold $\gamma_{th}=10$ dB. The mean and variance of UAV’s angular variations along the x-axis and y-axis are the same, with $\mu_{x}=\mu_{y}=0.5^{\circ}$ and $\sigma_{x}=\sigma_{y}=1^{\circ}$. To improve comprehension, we plot CDF instead of PDF for the array factor, single element radiation pattern, and radiation pattern. This transition from PDF to CDF provides a better understanding of the radiation patterns and their characteristics.
@@
Fig. 5 shows the CDF of the array factor, single element radiation pattern and radiation pattern of UMI for both Monte-Carlo simulations as well as analytical results calculated as $F_{\mathbb{G}_{a}}(\mathbb{G}_{a})=\sum_{i=1}^{J_{a}}p_{a}(i)U(\mathbb{G}_{a}-q_{a}(i))$, $F_{\mathbb{G}_{e}}(\mathbb{G}_{e})=\sum_{i=1}^{J_{e}}p(i)U(\mathbb{G}_{e}-q_{e}(i))$, and $F_{\mathbb{G}}(\mathbb{G})=\sum_{i=1}^{J}p(i)U(\mathbb{G}-q(i))$, respectively. Here, $U(x)=\begin{cases}1,&\mbox{}x\geq 0\\ 0,&\mbox{}x<0\end{cases}$. Furthermore, Fig. 5 illustrates that the accuracy of the analytical results is highly dependent on the number of sectors. A precise alignment between simulation and theoretical results can be achieved with a sufficiently large number of sectors. However, as shown in Fig. 5, even for smaller value of $N$, the analytical model obtained by using $D_{a}=15,D_{e}=5$ remains sufficiently accurate.
Figure 5: The CDF of (a) normalized array factor, (b) single element radiation pattern, and (c) radiation pattern of UMI.
@@
The performance of the UMI system is assessed using the OP as a key performance metric. To analyze the influence of the IRS pattern and the UAV’s angular vibrations on system performance, we compare scenarios without vibrations (marked as ’w/o’ in legends) against those with vibrations. Fig. 6 shows the outage probabilities of SISO- and MISO-Passive-IRS under CLT and Gamma assumptions, plotted as a function of $P_{t}$ for varying values of $N$. The simulation results confirm the accuracy of the derived analytical expressions. As shown in Fig. 6, the OP decreases with increasing the number of IRS elements $N$ when the UAV experiences no fluctuations. Conversely, when the UAV encounter a some fluctuations, the OP increases. From Fig. 6, it is clear that smaller values of $N$ result in improved performance in the high $P_{t}$ values, while larger values of $N$ improve performance at low $P_{t}$ values. This behavior can be attributed to the fact that at low $P_{t}$ values, stronger IRS directivity can effectively compensate for the signal. On the other hand, in high $P_{t}$ values, the OP for a narrow beam is constrained by the orientation fluctuations of the UAV. In such cases, a wider IRS pattern is necessary to mitigate the effects of UAV orientation changes. Moreover, the analytical outcome under the CLT and Gamma approximation closely aligns because they are equivalent for large $N$.
@@
Figure 6: The OP of SISO- and MISO-Passive-IRS systems versus transmit power, $P_{t}$ under CLT and Gamma assumption.
@@
Fig. 7 illustrates the OP of the SISO- and MISO-Active-IRS system as a function of transmit power, $P_{t}$ using CLT approximation for varying values of $N$. Based on our simulation results, we can confirm the accuracy of derived analytical expression. Consequently, the system utilizing active IRS employs fewer elements due to the amplification capability. As shown in Fig. 7, the system performance remains relatively stable with a low number of elements, even in the presence of vibrations. This stability can be attributed to the wider IRS pattern associated with fewer elements, which helps mitigate the impact of UAV fluctuations. Additionally, when the number of elements is low, the OP decreases with an increase in $N$ when UAV fluctuations are present. From Fig. 7, we can conclude that when UAVs experience fluctuations, employing active components with fewer elements is beneficial for minimizing the impact of these fluctuations. However, this approach may require additional power consumption.
@@
Figure 7: The OP of SISO- and MISO-Active-IRS system versus transmit power, $P_{t}$ under CLT approximation assumption.
@@
The optimal number of reflecting elements along with corresponding minimum achievable outage probabilities can be determined for varying values of $\mu_{x}$, $\mu_{y}$, $\sigma_{x}$, and $\sigma_{y}$. By analyzing the performance metrics based on these parameters, we can identify the configurations that yield the best reliability for the communication link in the UMI system. It is important to note that, based on our derived analytical expressions, the above optimization problem can be solved efficiently without requiring extensive computational time. Fig. 8 shows the OP for both passive and active IRS configurations in SISO and MISO systems as a function of the number of elements, $N$. In this analysis, the transmit power is set to $30$ dBm for the passive scenario and $5$ dBm for the active scenario, with $N_{\text{max}}=200$. As illustrated in Fig. 8, the optimal number of elements for each scenario is obtained using our derived analytical expressions. Note that the SISO-Passive-IRS configuration fails to operate effectively with transmit power equal to $30$ dBm, as the OP remains around $10^{-1}$ even with the maximum number of elements, $N=200$. In contrast, the MISO-Active-IRS system achieves a remarkably lower OP of $5\times 10^{-5}$ with optimal number of elements, $N_{\text{opt}}=50$. This significant difference highlights the advantages of using active components in the IRS, particularly in enhancing communication reliability under dynamic conditions.
Figure 8: The OP of all scenarios versus number of IRS elements, $N$: finding optimal number of IRS elements by solving optimization problem in (87).
@@ VI Conclusion
This study investigated the performance of an IRS mounted on a UAV, characterizing the IRS 3D pattern while accounting for random angular fluctuations of the UAV. We derived a closed-form PDF for the 3D IRS pattern to facilitate performance analysis and validated the accuracy of our analytical models using Monte Carlo simulations. The results demonstrated that, unlike terrestrial IRS systems, aerial IRS performance cannot be enhanced indefinitely by increasing the number of elements due to UAV instabilities. Additionally, simulations revealed that optimal system performance is achieved by selecting the optimal number of elements and incorporating active components.

### v3

@@ IV-B Active-IRS
$F_{\gamma_{max}|\mathbb{G}}(\gamma_{\text{th}})$ $=F_{U|Z}(\mu_{w}(\mathbb{G}))$ (75)
$+F_{U|Z}^{(1)}(\mu_{w}(\mathbb{G}))\mathbb{E}_{Z}\{(z-\mu_{z})\}$
$+F_{U|Z}^{(2)}(\mu_{w}(\mathbb{G}))\mathbb{E}_{Z}\{(z-\mu_{z})^{2}\}.$
Given that $\mathbb{E}_{Z}\{z-\mu_{z}\}=0$ and the third term in (75) is negligible compared to the first for a sufficient number of IRS elements $N$, we can approximate (75) as
$F_{\gamma_{max}|\mathbb{G}}(\gamma_{\text{th}})=Q(\frac{\mu_{v}-\sqrt{\mu_{w}(\mathbb{G})}}{\sigma_{v|z}})-Q(\frac{\mu_{v}+\sqrt{\mu_{w}(\mathbb{G})}}{\sigma_{v|z}}).$ (76)
Finally, substituting (76) into (68) yields the OP. ∎
Remark 4.
For a large number of active IRS elements ($N\rightarrow\infty$), the scaling $\mathbb{G}\propto 1/N^{2}$ is compensated by $U\propto N^{2}$ in the numerator of $\gamma_{\text{max}}=\frac{M\gamma_{0}\mathbb{G}U}{Z+c_{3}}$. Since the terms $Z$ and $c_{3}$ in the denominator scale proportionally with $N$, $\gamma_{\text{max}}\propto 1/N$. Consequently, the maximum SNR decays to zero under perfect CSI, forcing the OP to converge to $1$. This behavior yields a diversity order of zero, in contrast to the infinite diversity order achievable without fluctuations under perfect CSI. When CSI is imperfect, the OP for both scenarios converges to a constant floor dictated by the CSI error factor.
Remark 5.
For high SNR ($\gamma_{0}\rightarrow\infty$), the diversity order for both Passive-IRS and Active-IRS systems is limited and dictated by the UAV fluctuation level. This is in contrast to the no-fluctuations scenario, which achieves a diversity order of $N$ under perfect CSI [42]. When CSI is imperfect, the OP for both scenarios (with and without fluctuations) converges to a constant floor dictated by the CSI error factor, resulting in zero diversity order.
Remark 6.
The OP for both Passive-IRS and Active-IRS configurations yields the expression $P_{out}=\mathbf{p}^{T}T(\mathbf{q})$, where $T(\mathbf{q})$ denotes the element-wise application of function $T(\cdot)$ to vector $\mathbf{q}$. Computational evaluation requires: (i) generating vectors $\mathbf{p}$ and $\mathbf{q}$ through Kronecker products of component vectors of length $lD$, (ii) applying the transformation $T(\cdot)$ to each element of $\mathbf{q}$, and (iii) computing the dot product. Given that $J=l^{2}D^{2}$ is the dimension of the resulting vectors, each stage has complexity $\mathcal{O}(J)$, establishing the overall computational complexity as $\mathcal{O}(J)$.
@@ V Numerical results and discussions
This section provides numerical results to validate the theoretical analysis of the UMI system. The simulation parameters are configured as follows: the BS, IRS, and UE are located at $[0,0,20]$, $[10,10,120]$, and $[40,40,0]$, respectively. The number of BS antennas is $M=16$. The path-loss exponents for BS-IRS and IRS-UE links are configured as $\alpha_{1}=2$ and $\alpha_{2}=2.2$, respectively, with a reference path-loss of $c_{0}=-30$ dB applied to all links [43]. Additionally, both Rician factors for the BS-IRS and IRS-UE channels are set to $K_{0},K_{1}=10$ dB. The thermal noise power of the user’s receiver and active IRS is $\sigma_{n}^{2}=-80$ dBm and $\sigma_{f}^{2}=-70$ dBm, respectively [35]. The active IRS amplification power $P_{F}=0.05P_{t}$ and SNR threshold $\gamma_{th}=10$ dB. The mean and variance of UAV’s angular variations along the x-axis and y-axis are the same, with $\mu_{x}=\mu_{y}=0^{\circ}$ and $\sigma_{x}=\sigma_{y}=1^{\circ}$. To improve comprehension, we plot the CDF instead of the PDF for the UMI radiation pattern. This transition from PDF to CDF provides a better understanding of the radiation patterns and their characteristics.
@@
Fig. 4 shows the CDF of radiation pattern of UMI for both Monte-Carlo simulations as well as analytical results calculated as $F_{\mathbb{G}}(\mathbb{G})=\sum_{i=1}^{J}p(i)U(\mathbb{G}-q(i))$ where $U(x)=\begin{cases}1,&\mbox{}x\geq 0\\ 0,&\mbox{}x<0\end{cases}$. Furthermore, Fig. 4 illustrates that the accuracy of the analytical results is highly dependent on the number of sectors $D$ and the number of lobes $l$ (including both main- and side-lobes). A precise alignment between simulation and theoretical results can be achieved with a sufficiently large number of sectors. For instance, the results for $D=60$ ($1\%$ error) is markedly more accurate than for $D=15$ ($5\%$ error), at the cost of a $16$-fold increase in complexity, as noted in Remark 6. Furthermore, the analysis for $N=1024$ reveals that the narrow main lobe makes the main-lobe model ($l=1$) insufficient. Consequently, considering only the main-lobe ($l=1$) is insufficient. Accuracy is only achieved when the first side-lobe is included ($l=2$). In contrast, for other array sizes with wider main lobes, considering only the main lobe ($l=1$) provides an accurate approximation.
Figure 4: The CDF of the radiation pattern of UMI.
The performance of the UMI system is assessed using the OP as a key performance metric. To analyze the influence of the IRS pattern and the UAV’s angular vibrations on system performance, we compare scenarios without vibrations (marked as ’w/o vib’ in legends) against those with vibrations (marked as ’vib’ in legends). Our analysis reveals a direct trade-off between performance and computational cost governed by the number of sectors $D$. A larger $D$ yields a finer-grained approximation, which consistently improves the approximation accuracy (as measured by the computed OP), at the expense of increased computational complexity, which scales as $O(J)$ where $J=l^{2}D^{2}$ as derived in Remark 6.
@@
Fig. 5 shows the OP of the Passive-IRS under CLT and Gamma assumptions, plotted as a function of $P_{t}$ for varying values of the number of IRS elements $N$ and CSI error factor $\zeta$. The simulation results confirm the accuracy of the derived analytical expressions. As shown in Fig. 5, the OP decreases with increasing the number of IRS elements $N$ when the UAV experiences no fluctuations. Conversely, when the UAV encounters some fluctuations, the OP increases because of misalignment, as noted in Remark 1. From Fig. 5, it is clear that a smaller value of elements ($N=64$) results in improved performance in the high $P_{t}$ values, while a larger value of elements $N=256$ improves performance at low $P_{t}$ values. This behavior occurs because, at low $P_{t}$ values, the weaker IRS directivity (with $N=64$) cannot effectively compensate for the path loss. On the other hand, in high $P_{t}$ values, the OP for a narrow beam is constrained by the orientation fluctuations of the UAV, as noted in Remark 5. In such cases, a wider IRS pattern is necessary to mitigate the effects of UAV orientation changes. Moreover, the analytical outcome under the CLT and Gamma approximation closely aligns because they are equivalent for large $N$. The number of sectors $D$ in modeling the IRS radiation pattern impact on final results. Specifically, a larger $D=60$ ($2\%$ error) yields a more accurate model compared to a smaller $D=15$ ($8\%$ error). Furthermore, system performance degrades under imperfect CSI ($\zeta=0.1$) compared to the perfect CSI scenario ($\zeta=0$). This effect is more pronounced for a large number of elements, for instance $N=256$, because the variance of the CSI error is proportional to the cascaded channel gain, which itself scales with $N^{2}$.
@@
Figure 5: The OP of Passive-IRS systems versus transmit power, $P_{t}$ under CLT and Gamma approximation.
@@
Fig. 6 illustrates the OP of the Active-IRS system as a function of transmit power, $P_{t}$, using CLT approximation for varying values of the number of IRS elements $N$ and CSI error factor $\zeta$. Based on our simulation results, we can confirm the accuracy of the derived analytical expression. Consequently, the system utilizing active IRS employs fewer elements due to the amplification capability. As shown in Fig. 6, the system performance can be maintained relatively stable with a low number of elements, even in the presence of vibrations. This stability can be attributed to the wider IRS pattern associated with fewer elements, which helps mitigate the impact of UAV fluctuations. Additionally, for a given number of IRS elements, the Active-IRS system requires less transmit power than a Passive IRS to achieve the same performance. Alternatively, for a fixed transmit power, the Active-IRS system can achieve comparable performance with fewer elements. Since fewer elements have a wider pattern, this configuration is more robust to UAV fluctuations. From Fig. 6, we can conclude that when UAVs experience fluctuations, employing active components with fewer elements is beneficial for minimizing the impact of these fluctuations. However, this approach may require additional power consumption. The parameter $D$, which defines the number of sectors in modeling the IRS radiation pattern gain $\mathbb{G}$, impacts the final results. Specifically, a larger $D=60$ ($3\%$ error) yields a more accurate model compared to a smaller $D=15$ ($10\%$ error). Moreover, the system performance of Active-IRS, like Passive IRS, degrades under imperfect CSI ($\zeta=0.1$) compared to the perfect CSI scenario ($\zeta=0$). This effect is more pronounced for a large number of elements (e.g., $N=256$) because the CSI error variance scales with $N$ for active IRS but with $N^{2}$ for passive IRS, underscoring the superiority of the active IRS configuration under imperfect CSI.
@@
Figure 6: The OP of the Active-IRS system versus transmit power, $P_{t}$, under CLT approximation.
@@
The optimal number of reflecting elements along with corresponding minimum achievable outage probabilities can be determined for varying values of $\mu_{x}$, $\mu_{y}$, $\sigma_{x}$, and $\sigma_{y}$. By analyzing the performance metrics based on these parameters, we can identify the configurations that yield the best reliability for the communication link in the UMI system. It is important to note that, based on our derived analytical expressions, the above optimization problem can be solved efficiently without requiring extensive computational time.
Fig. 7 shows the OP for both passive and active IRS configurations under perfect ($\zeta=0$) and imperfect ($\zeta=0.1$) CSI as a function of the number of elements, $N$. In this analysis, the transmit power is set to $30$ dBm for the passive scenario and 0 dBm for the active scenario, with $N_{\text{max}}=400$. As illustrated in Fig. 7, the optimal number of IRS elements for each scenario is determined using our derived analytical expressions. The results show optimal values of $N_{\text{opt}}=144$ for the Passive-IRS and $N_{\text{opt}}=49$ for the Active-IRS system. While the Passive-IRS with its optimal configuration achieves the OP of $8\times 10^{-3}$ under imperfect CSI, the Active-IRS achieves a remarkably lower OP of $1\times 10^{-4}$. This significant difference highlights the advantages of using active components in the IRS, particularly in enhancing communication reliability under dynamic conditions. Furthermore, Fig. 7 reveals that the value of $N_{\text{opt}}$ is identical for $D=15$ and $D=30$, emphasizing that a finer sectorization beyond $D=15$ is not necessary for this specific optimization. Additionally, while imperfect CSI ($\zeta=0.1$) leads to a higher overall OP, it does not affect the optimal number of elements; it only shifts the OP performance curve. As shown in Fig. 7, the diversity order for both IRS systems is zero in the presence of UAV fluctuations, whereas an infinite diversity order is achieved in the absence of fluctuations, even with perfect CSI, as analytically noted in Remarks 3 and 4.
Figure 7: The OP of all scenarios versus the number of IRS elements, NN: finding the optimal number of IRS elements by solving the optimization problem in (77).
@@ VI Conclusion
This study investigated the performance of an IRS mounted on a UAV, characterizing the IRS 3D pattern while accounting for random angular fluctuations of the UAV. We derived a closed-form PDF for the 3D IRS pattern to facilitate performance analysis and validated the accuracy of our analytical models using Monte Carlo simulations. The results demonstrated that, unlike terrestrial IRS systems, aerial IRS performance cannot be enhanced indefinitely by increasing the number of elements due to UAV instabilities. Additionally, simulations revealed that optimal system performance is achieved by selecting the optimal number of elements and incorporating active components. Building upon the foundational understanding established in this work, our future research will extend this framework to a multi-user scenario.